# DIAS: A Comprehensive Dataset and Benchmark for Intracranial Artery Segmentation in DSA sequences

### 12 Dec 2023 v2

@@ V-C Weakly-supervised Segmentation
We compared our method with five scribble-supervised segmentation methods: partial Cross Entropy (pCE) [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40]. To ensure a fair comparison, all experimental networks employed the UNet [27], consistent with the approach in DMPLS. Table III presents the results of scribble-supervised segmentation of the DIAS dataset. Fig. 7 showcases the visual results of all methods. Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods. It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 4.67% higher compared to SSCR, and it is 4.12% higher on SALE. Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background. The visualization results from IVM indicate that achieving high Sen comes at the cost of introducing numerous incorrectly segmented vessel pixels. Consequently, this leads to lower scores in other evaluation metrics such as the DSC and AUC. Regarding SALE, the vessels segmented by the same method are thicker than the ones trained with RDFA annotation, leading to a larger number of false positive pixels at the vessel’s edge. This outcome may have been caused by the SALE focusing on the annotation of thick vessels. Similarly, due to the lack of supervision for small vessels, where vascular connectivity issues are more prevalent, SALE struggles to extract fine vessels, resulting in a VC score significantly better than RDFA. From the last column of the figure, it is evident that SSCR’s segmentation results have the least number of blue pixels, corroborating the best quantitative results presented in the Table III.
@@
Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.591 0.9694 148.89
Student 10 30 0.7544 0.9618 0.766 0.9786 0.6087 0.9769 93.31
Student w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
Student 10 60 0.7579 0.9625 0.7662 0.9795 0.613 0.9786 87.72
We conducted ablation studies to investigate the impact of various components in our proposed pipeline for weakly-supervised IA segmentation. As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis. Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as DSC and AUC. Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC. Compared to a single-model approach, a dual-model setup, receiving the same scribble supervision, generates pseudo-labels for mutual supervision. This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability. We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation without CPS. Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC. Additionally, it has attained the highest Sen. We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models. The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 2.5% and 2.38%, respectively.
@@ VI-B Potential Clinical Applications and Limitations
Segmentation of vessels in DSA for quantifying their morphology has already found application in clinical research, particularly for planning interventional surgeries [7] and diagnosing strokes [5]. However, as discussed in Section VI-A, the segmentation of a single DSA frame is inadequate for fully delineating the vascular structure. We believe that the application of our proposed method for segmenting DSA vessel sequences in the aforementioned higher-level tasks will yield more accurate vascular structures. This advancement is expected to drive forward the progress of these higher-level tasks. In this work, we propose a comprehensive benchmark based on our created dataset for intracranial arterial segmentation in DSA sequences. Importantly, the proposed approach is not confined to intracranial arteries. With the availability of relevant annotated data, it can be extended to other segmentation tasks in DSA sequences, such as venous segmentation [8], as well as for the diagnosis of stroke-related conditions like aneurysms [3], ICAS, and MCAO [59].
Currently, the DIAS dataset is relatively small, primarily due to the substantial effort expended on one-to-one pixel annotations of DSA sequences. In future work, we plan to leverage RPST to generate pseudo-labels for DSA sequences pending annotation. Annotation experts will then meticulously revise these pseudo-labels to obtain precise annotations, rather than starting the annotation process from the beginning. This strategy aims to increase annotation efficiency, thereby facilitating the further expansion of DIAS. On the other hand, in the DIAS dataset, intracranial arterial sequences are selected by neurosurgeons from DSA sequences that may include stages such as the capillary or venous phase. Given the nature of 3D spatial projection, angiography from other phases in DSA may impact arterial segmentation, as the same pixel can represent an artery or a vein at different times. In our next phase of work, we plan to segment intracranial arteries from DSA sequences that include the venous phase and other stages, without relying on the neurosurgeons’ prior knowledge to pre-select arterial phase sequences.
@@ VII Conclusion
Acknowledgments
This work was supported in part by the National Natural Science Foundation of China (No.62376038) and National Key R&D Program of China (No.2018AAA0102600).

### 17 Dec 2023 v3

@@ V-C Weakly-supervised Segmentation
We compared our method with five scribble-supervised segmentation methods: pCE [57], Entropy Minimization (EM) [56], Regularized Loss (RLoss) [45], Intra-class Intensity Variance Minimization (IVM) [46] and Dynamically Mixed Pseudo Labels Supervision (DMPLS) [40]. To ensure a fair comparison, all experimental networks employed the UNet [27], consistent with the approach in DMPLS. Table III presents the results of scribble-supervised segmentation of the DIAS dataset. Fig. 7 showcases the visual results of all methods. Thanks to pseudo supervision and consistency regularization for scribble supervision, SSCR achieved significantly better performance in RDFA and SALE by obtaining the highest scores in most metrics among all methods. It is noteworthy that in the segmentation results of IVM, the Sen on RDFA is 4.67% higher compared to SSCR, and it is 4.12% higher on SALE. Sen is used to measure the proportion of correctly identified positive classes (i.e., vessels) out of all actual positive classes, independent of the background. The visualization results from IVM indicate that achieving high Sen comes at the cost of introducing numerous incorrectly segmented vessel pixels. Consequently, this leads to lower scores in other evaluation metrics such as the DSC and AUC. Regarding SALE, the vessels segmented by the same method are thicker than the ones trained with RDFA annotation, leading to a larger number of false positive pixels at the vessel’s edge. This outcome may have been caused by the SALE focusing on the annotation of thick vessels. Similarly, due to the lack of supervision for small vessels, where vascular connectivity issues are more prevalent, SALE struggles to extract fine vessels, resulting in a VC score significantly better than RDFA. From the last column of the figure, it is evident that SSCR’s segmentation results have the least number of blue pixels, corroborating the best quantitative results presented in the Table III.
@@
Teacher 10 0 0.7402 0.9596 0.7477 0.9779 0.5910 0.9694 148.89
Student 10 30 0.7544 0.9618 0.7660 0.9786 0.6087 0.9769 93.31
S w/o SDA 10 60 0.7556 0.9625 0.7552 0.9805 0.6096 0.9737 93.92
Student 10 60 0.7579 0.9625 0.7662 0.9795 0.6130 0.9786 87.72
We conducted ablation studies to investigate the impact of various components in our proposed framework for weakly-supervised IA segmentation. As illustrated in Table IV, with pCE serving as the baseline, we incrementally added components for comparative analysis. Initially, the introduction of pseudo-label supervision resulted in a slight decline in metrics such as DSC and AUC. Based on this, we utilized a dual-model framework to establish cross pseudo-supervision, leading to a 1.08% enhancement in DSC. Compared to a single-model approach, a dual-model setup, receiving the same scribble supervision, generates pseudo-labels for mutual supervision. This avoids the scenario of a model being self-supervised by its own predictions, thereby enhancing the model’s generalizability. We employ consistency regularization on the pseudo-labels generated by dual models to penalize inconsistent segmentation. Compared to ‘pCE+PS’, ‘pCE+PS+CR’ exhibits a 1.2% improvement in DSC. Additionally, it has attained the highest Sen. We implement a combined strategy of cross pseudo-supervision and consistency regularization on the pseudo-labels produced by dual models. The integrated approach, termed as SSCR and denoted as ‘pCE+PS+CPS+CR’, has achieved notable improvements, surpassing ‘pCE+PS+CPS’ and ‘pCE+PS+CR’ by 2.5% and 2.38%, respectively.
@@ VI-B Potential Clinical Applications and Limitations
Segmentation of vessels in DSA for quantifying their morphology has already found application in clinical research, particularly for planning interventional surgeries [7] and diagnosing strokes [5]. However, as discussed in Section VI-A, the segmentation of a single DSA frame is inadequate for fully delineating the vascular structure. We believe that the application of VSS-Net for segmenting DSA vessel sequences in the aforementioned higher-level tasks will yield more accurate vascular structures. This advancement is expected to drive forward the progress of these higher-level tasks. Although the SSCR using scribble annotations falls short compared to fully supervised segmentation, the results shown in Fig. 7 reveal that this shortfall is primarily in the segmentation of smaller, terminal branches of vessels. Conversely, the segmentation of main vessels such as ICA, ACA, and MCA is comparatively effective. This is of clinical importance, as stroke-related ailments such as aneurysms [3], ICAS, and MCAO [59] predominantly occur in these main vessels. It is noteworthy that our assessment indicates the time taken for scribble annotation of a DSA sequence is just one-tenth of that required for full annotation. In this work, we introduce the FSS, WSS, and SSS methods for intracranial artery segmentation in DSA sequences, specifically designed for the DIAS dataset. Importantly, the proposed approach is not confined to intracranial arteries. With the availability of relevant annotated data, it can be extended to other segmentation tasks in DSA sequences, such as venous segmentation [8], as well as for the diagnosis of stroke-related conditions like aneurysms, ICAS, and MCAO.
The experimental results of the RPST demonstrates the potential of semi-supervised segmentation methods in the segmentation of IA in DSA sequences. One straightforward application is in assisting with data annotation. Currently, the DIAS dataset is relatively small, primarily due to the substantial effort expended on one-to-one pixel annotations of DSA sequences. In future work, we plan to leverage RPST to generate pseudo-labels for DSA sequences pending annotation. Annotation experts will then meticulously revise these pseudo-labels to obtain precise annotations, rather than starting the annotation process from the beginning. This strategy aims to increase annotation efficiency, thereby facilitating the further expansion of DIAS. On the other hand, in the DIAS dataset, intracranial arterial sequences are selected by neurosurgeons from DSA sequences that may include stages such as the capillary or venous phase. Given the nature of 3D spatial projection, angiography from other phases in DSA may impact arterial segmentation, as the same pixel can represent an artery or a vein at different times. In our next phase of work, we plan to segment intracranial arteries from DSA sequences that include the venous phase and other stages, without relying on the neurosurgeons’ prior knowledge to pre-select arterial phase sequences.